OpenStudy (anonymous):

FInd 2 posetive numbers such that their product is 185 and their sum is a minimum. (optimization)

hartnn (hartnn):

using calculus ?

OpenStudy (anonymous):

Yes :)

hartnn (hartnn):

let the 2 numbers be x and y so , xy= ? function to be minimised = x+y= ?

OpenStudy (anonymous):

xy=185 i dont understand the 2nd part

hartnn (hartnn):

xy =185 so find y from this(or x) then put it in f=x+y

OpenStudy (anonymous):

oh. so do i put f in for y in xy=185?

hartnn (hartnn):

y=185/x f(x) = x + 185/x got this ? now miminize f(x) know how to ?

OpenStudy (anonymous):

ohhhh! ok. let me process this x)

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

Take the derivative?

hartnn (hartnn):

yes! take derivative and equate it to ..... ?

OpenStudy (anonymous):

0 ~(^_^)~

hartnn (hartnn):

absolutely!

hartnn (hartnn):

so what derivative u got ? f'(x) = ?

OpenStudy (anonymous):

1+ (185/x^2)

OpenStudy (anonymous):

then do I solve for x?

hartnn (hartnn):

sure? no sign mistakes ?

OpenStudy (anonymous):

Oops! Its "-" not "+" x)

hartnn (hartnn):

right. and yes now equate that to 0

hartnn (hartnn):

1- (185/x^2) =0 x= ?

OpenStudy (anonymous):

\[\sqrt{185}\]

hartnn (hartnn):

yes! actually u also get -sqrt(185) but since the numbers are positive, u only get x=+sqrt(185) got this ?

OpenStudy (anonymous):

oh, yeah. Got it.

OpenStudy (anonymous):

Now do I take the second derivitive?

hartnn (hartnn):

i don't think that required, u got x, now find y and u are done.....

OpenStudy (anonymous):

So.. find y by putting the answere I got for x into the xy=185?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

the answere is \[\sqrt{185}\] for both! :D

hartnn (hartnn):

yes!

OpenStudy (anonymous):

Thank u sooo much!

hartnn (hartnn):

welcome sooo much!